# Complex Domain Approach for Reversible Data Hiding and Homomorphic Encryption: General Framework and Application to Dispersed Data

### v1

@@ V-A1 Complexity analysis
For a Gaussian integer $\chi=x+iy\in{\mathbb{Z}[i]}^{*}_{p}$, its modular inverse $\chi^{-1}\,(\operatorname{\mathrm{mod}}p)$ is typically computed as $\overline{\chi}\cdot(\mathrm{N}(\chi))^{-1}\,(\operatorname{\mathrm{mod}}p)$. This involves obtaining the modular inverse of the integer norm $\mathrm{N}(\chi)=x^{2}+y^{2}$ modulo $p$ (an integer modular exponentiation), followed by a complex multiplication.
@@
Regarding the relative computational cost, a complex modular exponentiation is approximately four times as expensive as an integer modular exponentiation of a similar exponent length. This is primarily because each complex modular multiplication, which forms the building block of exponentiation, involves four integer modular multiplications. $\triangle$
The computational overhead for each party during a single data exchange round is analyzed in terms of equivalent integer modular exponentiations (as defined in Remark 10). The Sensor performs approximately 8 equivalent integer modular exponentiations for encrypting its watermarked data. In contrast, the Data Collector incurs a higher computational burden, totaling approximately 13 equivalent integer modular exponentiations per round (and sensor). The majority of this computational load at the DC is primarily due to the operations involved in encrypting the challenge factor ($\lambda_{k}$) sent to the Sensor, decrypting the combined ciphertext received from the Sensor, and then isolating the original watermarked data. This distribution of workload, where the Sensor’s load is comparatively lighter, aligns well with typical sensor network architectures where sensors are often resource-constrained devices, while the data collector usually possesses greater computational power.
V-B H[i][i]dden-AggP: Joint Watermarking, Encryption, and Aggregation
@@
1) Paillier Public Key Distribution: DC generates a Paillier key pair. It computes and publicly distributes its public key $\mathsf{K_{pub}}=(n,g)$, where $n=pq$ for large primes $p,q$, and $g$ is a generator as defined in Section II-B2. DC keeps its corresponding private key $\mathsf{K_{priv}}=(L_{P},M_{P})$ secret.
2) Symmetric Key Exchange (for DC to Sensor): For each sensor $\mathrm{S}_{j}$ ($j=1,\ldots,N$), $\mathrm{S}_{j}$ generates a unique, random symmetric key $K_{\mathrm{sym},j}$ (e.g., an AES key). $\mathrm{S}_{j}$ then encrypts $K_{\mathrm{sym},j}$ using DC’s public Paillier key $\mathsf{K_{pub}}$, resulting in ciphertext $c_{j}=\mathrm{Enc}_{\mathsf{K_{pub}}}(K_{\mathrm{sym},j},r_{j})$, where $r_{j}$ is a random integer. $\mathrm{S}_{j}$ transmits $c_{j}$ to DC. Upon reception, DC decrypts $c_{j}$ using its private key $\mathsf{K_{priv}}$ to obtain $K_{\mathrm{sym},j}$. This process establishes a unique symmetric key $K_{\mathrm{sym},j}$ shared securely between DC and each individual sensor $\mathrm{S}_{j}$. These symmetric keys will subsequently be used for secure communication from DC to $\mathrm{S}_{j}$.
3) Watermark Agreement: DC and all sensors $\mathrm{S}_{j}$ collectively agree on a set of shared, sequential, time-dependent watermarks: $w_{1},w_{2},\ldots,w_{M}$ corresponding to times $t_{1},t_{2},\ldots,t_{M}$. Each $w_{k}$ (for the $k$-th round) is a non-negative integer represented with $B$ bits, taking values in the range $0,\ldots,2^{B}-1$.
@@
1) DC’s Challenge Factor Distribution: DC chooses a Gaussian integer factor $\lambda_{k}$ with $\operatorname{Re}(\lambda_{k})\cdot\operatorname{Im}(\lambda_{k})\neq 0$. The real and imaginary parts of $\lambda_{k}$ must be chosen such that the resulting components of $\delta^{\prime}_{j,k}$ can be represented within Paillier’s plaintext modulus $n$. For each sensor $\mathrm{S}_{j}$ ($j=1,\ldots,N$), DC encrypts $\lambda_{k}$ using the pre-established unique symmetric key $K_{\mathrm{sym},j}$ shared with that specific sensor. Let this symmetric encryption be $E_{K_{\mathrm{sym},j}}(\lambda_{k})$ (implying $\lambda_{k}$ is suitably serialized for symmetric encryption). DC then sends $E_{K_{\mathrm{sym},j}}(\lambda_{k})$ to each $\mathrm{S}_{j}$. Each sensor $\mathrm{S}_{j}$ decrypts the received message using its $K_{\mathrm{sym},j}$ to obtain the plaintext $\lambda_{k}$.
2) Sensor Data Embedding and Initial Paillier Encryption: Each sensor $\mathrm{S}_{j}$ generates its data reading $d_{j,k}$ (an integer). $\mathrm{S}_{j}$ then combines the pre-agreed watermark $w_{k}$ with its data, forming a complex number $\delta_{j,k}=d_{j,k}+iw_{k}$. Using the received $\lambda_{k}$, $\mathrm{S}_{j}$ then computes the scaled and watermarked complex number $\delta^{\prime}_{j,k}=\lambda_{k}\delta_{j,k}$. This $\delta^{\prime}_{j,k}$ is the value that will be aggregated. $\mathrm{S}_{j}$ encrypts $\delta^{\prime}_{j,k}$ component-wise using DC’s public Paillier key $\mathsf{K_{pub}}$. Let $\delta^{\prime}_{j,k}=m_{j,k,R}+im_{j,k,I}$ (where $m_{j,k,R}$ and $m_{j,k,I}$ are integers). The ciphertext is $(c_{j,k,R},c_{j,k,I})=(\mathrm{Enc}_{\mathsf{K_{pub}}}(m_{j,k,R},r_{j,k,R}),\mathrm{Enc}_{\mathsf{K_{pub}}}(m_{j,k,I},r_{j,k,I}))$, where $r_{j,k,R}$ and $r_{j,k,I}$ are fresh random numbers for each encryption.
3) Cascading Encrypted Summation (Tree Aggregation): The individual encrypted watermarked data points are then aggregated in a cascading manner, typically following a binary tree structure involving $\lceil\log_{2}(N)\rceil$ steps.
@@
  • Aggregation Steps: For each internal node in the aggregation tree, a designated sensor (or an intermediate aggregator) receives two ciphertexts, say $(C_{1R},C_{1I})$ and $(C_{2R},C_{2I})$, from its children (or from other sensors). Taking advantage of the additive homomorphic property of Paillier encryption, this aggregator computes the sum of the encrypted real parts ($C_{\mathrm{sum},R}=C_{1R}\cdot C_{2R}\,(\operatorname{\mathrm{mod}}n^{2})$) and the sum of the encrypted imaginary parts ($C_{\mathrm{sum},I}=C_{1I}\cdot C_{2I}\,(\operatorname{\mathrm{mod}}n^{2})$). The resulting ciphertext $(C_{\mathrm{sum},R},C_{\mathrm{sum},I})$ is then passed up the tree. This process continues until a single final encrypted sum is obtained.
@@
The final encrypted sum, $\sigma^{\prime}_{\mathrm{sum},k}=(C_{\mathrm{final},R},C_{\mathrm{final},I})$, is obtained by the root sensor in the aggregation tree (e.g., $\mathrm{S}_{N}$ if sensors are ordered by aggregation path) and is then sent to DC.
@@
The decrypted value is the aggregated watermarked data $\sigma^{\prime}_{k}=\sigma^{\prime}_{k,R}+i\sigma^{\prime}_{k,I}$. As derived in Section III-C, $\sigma^{\prime}_{k}=\lambda_{k}(S_{k}+iNw_{k})$, where $S_{k}=\sum_{j=1}^{N}d_{j,k}$. DC then isolates the true aggregated data $S_{k}$ and the watermark $w_{k}$ by computing $(S_{k}+iNw_{k})=\sigma^{\prime}_{k}/\lambda_{k}$ using standard complex number division. Finally, DC extracts $S_{k}$ from the real part and obtains $w_{k}$ by dividing the imaginary part by $N$. If the extracted watermark $w_{k}$ correctly matches the expected time-dependent watermark, the aggregated data $S_{k}$ is accepted; otherwise, it is discarded, indicating potential tampering or an unauthorized source.
V-B1 Complexity Analysis
@@
  • DC to Sensors ($\lambda_{k}$ distribution): The DC sends the symmetrically encrypted challenge factor $E_{K_{\mathrm{sym},j}}(\lambda_{k})$ to each of the $N$ sensors. This accounts for $N$ messages from DC to sensors.

### v2

@@ 5.2.1 Complexity analysis
For a Gaussian integer $\chi=x+iy\in{\mathbb{Z}[i]}^{*}_{p}$, its modular inverse $\chi^{-1}\,(\operatorname{\mathrm{mod}}p)$ is typically computed as $\overline{\chi}\cdot(\mathrm{N}(\chi))^{-1}\,(\operatorname{\mathrm{mod}}p)$. This involves obtaining the modular inverse of the integer norm $\mathrm{N}(\chi)=x^{2}+y^{2}$ modulo $p$ (an integer modular exponentiation), followed by a complex multiplication (Fermat’s Little Theorem):
$(x^{2}+y^{2})^{-1}\,(\operatorname{\mathrm{mod}}p)=(x^{2}+y^{2})^{p-2}\,(\operatorname{\mathrm{mod}}p).$
@@
Regarding the relative computational cost, a complex modular exponentiation is approximately four times as expensive as an integer modular exponentiation of a similar exponent length. This is primarily because each complex modular multiplication, which forms the building block of exponentiation, involves four integer modular multiplications:
$(a+ib)(c+id)\,(\operatorname{\mathrm{mod}}p)=((ac-bd)+i(ad+bc))\,(\operatorname{\mathrm{mod}}p).$
$\triangle$
The computational cost for each entity during a single data exchange round is analyzed in terms of equivalent integer modular exponentiations (as defined in Remark 15). The Sensor performs approximately 8 equivalent integer modular exponentiations for encrypting its watermarked data. In contrast, the Data Collector incurs a higher computational burden, totaling approximately 13 equivalent integer modular exponentiations per round (and sensor). The majority of this computational load at the DC is primarily due to the operations involved in encrypting the challenge factor ($\lambda_{k}$) sent to the Sensor, decrypting the combined ciphertext received from the Sensor, and then isolating the original watermarked data. This distribution of workload, where the Sensor’s load is comparatively lighter, aligns well with typical sensor network architectures where sensors are often resource-constrained devices, while the data collector usually possesses greater computational power.
Figure 2: Graphical representation of the H[i][i]dden-AggP protocol, including the initialization (above the horizontal dashed line) and the data exchange (below the horizontal dashed line) phases.
5.3 H[i][i]dden-AggP: Joint Watermarking, Encryption, and Aggregation
@@
1) Paillier Public Key Distribution: DC generates a Paillier key pair. It computes and publicly distributes its public key $\mathsf{K_{pub}}=(n,g)$, where $n=pq$ for large primes $p,q$, and $g$ is a generator as defined in Section 2.2.2. DC keeps its corresponding private key $\mathsf{K_{priv}}=(L_{P},M_{P})$ secret.
2) Symmetric Key Exchange (for DC to Sensor): Each sensor $\mathrm{S}_{j}$ ($j=1,\ldots,N$), $\mathrm{S}_{j}$ generates a unique, random symmetric key $K_{\mathrm{sym},j}$ (e.g., an AES key). $\mathrm{S}_{j}$ then encrypts $K_{\mathrm{sym},j}$ using DC’s public Paillier key $\mathsf{K_{pub}}$, resulting in ciphertext $c_{j}=\mathrm{Enc}_{\mathsf{K_{pub}}}(K_{\mathrm{sym},j},r_{j})$, where $r_{j}$ is a random integer. $\mathrm{S}_{j}$ transmits $c_{j}$ to DC. Upon reception, DC decrypts $c_{j}$ using its private key $\mathsf{K_{priv}}$ to obtain $K_{\mathrm{sym},j}$. This process establishes a unique symmetric key $K_{\mathrm{sym},j}$ shared securely between DC and each individual sensor $\mathrm{S}_{j}$. These symmetric keys will subsequently be used for secure communication from DC to $\mathrm{S}_{j}$.
3) Watermark Agreement: DC and all sensors $\mathrm{S}_{j}$ collectively agree on a set of shared, sequential, time-dependent watermarks: $w_{1},w_{2},\ldots,w_{M}$ corresponding to times $t_{1},t_{2},\ldots,t_{M}$. Each $w_{k}$ (for the $k$-th round) is an integer represented with $B$ bits, which may be signed or unsigned as detailed in the H[i][i]dden-EG case.
@@
1) DC’s Challenge Factor Distribution: DC chooses a Gaussian integer factor $\lambda_{k}$ with $\operatorname{Re}(\lambda_{k})\cdot\operatorname{Im}(\lambda_{k})\neq 0$. The real and imaginary parts of $\lambda_{k}$ must be chosen such that the resulting components of $\delta^{\prime}_{j,k}$ can be represented within Paillier’s plaintext modulus $n$. For each sensor $\mathrm{S}_{j}$ ($j=1,\ldots,N$), DC encrypts $\lambda_{k}$ using the pre-established unique symmetric key $K_{\mathrm{sym},j}$ shared with that specific sensor. Let this symmetric encryption be $\mathrm{Enc}_{K_{\mathrm{sym},j}}(\lambda_{k})$ (implying $\lambda_{k}$ is suitably serialized for symmetric encryption). DC then sends $\mathrm{Enc}_{K_{\mathrm{sym},j}}(\lambda_{k})$ to each $\mathrm{S}_{j}$. Each sensor $\mathrm{S}_{j}$ decrypts the received message using its $K_{\mathrm{sym},j}$ to obtain the plaintext $\lambda_{k}$.
2) Sensor Data Embedding and Initial Paillier Encryption: Each sensor $\mathrm{S}_{j}$ generates its data reading $d_{j,k}$ (an integer). $\mathrm{S}_{j}$ then combines the pre-agreed watermark $w_{k}$ with its data, forming a complex number $\delta_{j,k}=d_{j,k}+iw_{k}$. Using the received $\lambda_{k}$ (once decrypted), $\mathrm{S}_{j}$ then computes the scaled and watermarked complex number $\delta^{\prime}_{j,k}=\lambda_{k}\delta_{j,k}$. This $\delta^{\prime}_{j,k}$ is the value that will be aggregated. $\mathrm{S}_{j}$ encrypts $\delta^{\prime}_{j,k}$ component-wise using DC’s public Paillier key $\mathsf{K_{pub}}$. Let $\delta^{\prime}_{j,k}=m_{j,k,R}+im_{j,k,I}$ (where $m_{j,k,R}$ and $m_{j,k,I}$ are integers). The ciphertext is $(c_{j,k,R},c_{j,k,I})=(\mathrm{Enc}_{\mathsf{K_{pub}}}(m_{j,k,R},r_{j,k,R}),\mathrm{Enc}_{\mathsf{K_{pub}}}(m_{j,k,I},r_{j,k,I}))$, where $r_{j,k,R}$ and $r_{j,k,I}$ are fresh random numbers for each encryption.
3) Cascading Encrypted Summation (Tree Aggregation): The individual encrypted watermarked data points are then aggregated in a cascading manner, typically following a binary tree structure involving $\lceil\log_{2}(N)\rceil$ steps:
@@
  • Aggregation Steps: For each internal node in the aggregation tree, a designated sensor (or an intermediate aggregator) receives two ciphertexts, say $(C_{1R},C_{1I})=({c_{j_{1},k,R},c_{j_{1},k,I})}$ and $(C_{2R},C_{2I})=({c_{j_{2},k,R},c_{j_{2},k,I})}$, from its children (or from other sensors). Taking advantage of the additive homomorphic property of Paillier encryption, this aggregator computes the sum of the encrypted real parts ($C_{\mathrm{sum},R}=C_{1R}\cdot C_{2R}\,(\operatorname{\mathrm{mod}}n^{2})$) and the sum of the encrypted imaginary parts ($C_{\mathrm{sum},I}=C_{1I}\cdot C_{2I}\,(\operatorname{\mathrm{mod}}n^{2})$). The resulting ciphertext $(C_{\mathrm{sum},R},C_{\mathrm{sum},I})$ is then passed up the tree. This process continues until a single final encrypted sum is obtained.
@@
The final encrypted sum, $\sigma^{\prime}_{\mathrm{sum},k}=(C_{\mathrm{final},R},C_{\mathrm{final},I})$, is obtained by the root sensor in the aggregation tree (e.g., $\mathrm{S}_{1}$ if sensors are ordered by aggregation path) and is then sent to DC.
@@
The decrypted value is the aggregated watermarked data $\sigma^{\prime}_{k}=\sigma^{\prime}_{k,R}+i\sigma^{\prime}_{k,I}$. As derived in Section 3.3, $\sigma^{\prime}_{k}=\lambda_{k}(S_{k}+iNw_{k})$, where $S_{k}=\sum_{j=1}^{N}d_{j,k}$. DC then isolates the true aggregated data $S_{k}$ and the watermark $w_{k}$ by computing $(S_{k}+iNw_{k})=\sigma^{\prime}_{k}/\lambda_{k}$ using standard complex number division. Finally, DC extracts $S_{k}$ from the real part and obtains $w_{k}$ by dividing the imaginary part by $N$. If the extracted watermark $w_{k}$ correctly matches the expected time-dependent watermark, the aggregated data $S_{k}$ is accepted; otherwise, it is discarded, indicating potential tampering or an unauthorized source.
Remark 16.
The component-wise application of Paillier encryption requires the real and imaginary parts of the aggregate $\mu=\lambda_{k}(S_{k}+iNw_{k})$ to be handled modulo $n$. In this scenario, the additive homomorphic property allows for the accumulation of $N$ sensor readings. To ensure that the aggregate result does not exceed the capacity of the modulus $n$ and avoids modular wrap-around, the parameters must satisfy:
$\max(|\operatorname{Re}(\mu)|,|\operatorname{Im}(\mu)|)<n/2.$
This condition ensures that both components of the aggregate Gaussian integer can be correctly mapped back from the centered representative system of $\mathbb{Z}_{n}$ to their exact integer values in $\mathbb{Z}$, which is a prerequisite for the successful application of the perfect reversibility proven in Theorem 6. $\triangle$
Figure 3: Cascading summation protocol: example for 11 sensors.
A graphical representation of the H[i][i]dden-AggP protocol is shown in Fig. 2, whereas Fig. 3 illustrates how the cascading summation would work for 11 sensors. In this figure, the $\mathrm{S}_{1}$ sensor would obtain the final aggregate result to send it to the DC. It must be taken into account that the $\Sigma$” symbols in the figure represent multiplication of ciphertexts for the Paillier cryptosystem.
Remark 17.
The cascading aggregation protocol shown in Fig. 3 offers the advantage of performing the necessary summation internally within the sensor network without requiring a dedicated external entity. While this approach distributes the computational load, the topology implies a sequential delay of $\lceil\log_{2}(N)\rceil$ communication steps as partial aggregates propagate. Alternatively, if a participating sensor (e.g., $\mathrm{S}_{1}$) is sufficiently equipped with storage and computational power, it can act as a hub, receiving all ciphertexts directly from other nodes. While this reduces the transmission latency to a single step, the node must still perform N1N-1 modular multiplications (modulo $n^{2}$) sequentially to produce the final result. In fact, whether using a tree or a star arrangement, the algebraic work remains consistent. Performing aggregation within the sensor network itself rather than employing an external Data Aggregator (DA) minimizes the attack surface and circumvents the complex security considerations of third-party collusion (e.g., potential collusions between the DA and the DC). $\triangle$
5.3.1 Complexity Analysis
@@
  • DC to Sensors ($\lambda_{k}$ distribution): The DC sends the symmetrically encrypted challenge factor $\mathrm{Enc}_{K_{\mathrm{sym},j}}(\lambda_{k})$ to each of the $N$ sensors. This accounts for $N$ messages from DC to sensors.